kekeman:

The population density of Appleland is 12 apple trees per acre. Exactly 792 apple trees grow in Appleland. How many acres are in Appleland? 40 44 53 66

thony:

If there are 12 trees per acre then 792/12

jhonyy9:

just use the rule of 3 simple and write in this way 12 trees per 1 acre 792 trees per x acres and cross multiplie to calcule the x value x = 792*1/12 = 792/12 = ? acres

kekeman:

66

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kekeman 66 \(\color{#0cbb34}{\text{End of Quote}}\) yes ofc